In a Math contest, Team B earned 3 points less than Team C. Team A garnered 6 points more than Team B. What is the difference between the highest and lowest score?
Answer by stanbon(48558)

In a Math contest, Team B earned 3 points less than Team C. Team A garnered 6 points more than Team B. What is the difference between the highest and lowest score?
------------
If Team C has "x" points,
Team B has "x-3" points.
---------
The Team A has (x-3)+6 = x+3 points
---------
The lowest is "x-3"
The highest is "x+3"
-----------------
highest - lowest = x+3-(x-3) = 6 points
===============
Cheers,
Stan H.